中級技術英語 1 単位 (選択)

Technical English  カーペンターウォルター · 特任講師 / 機械工学科生産システム講座

【授業目的】 The aim of the course is to enhance reading, writing, speaking and listening skills. Another goal is to establish a firm foundation for students to present their knowledge to others, in English.

【授業概要】 The primary focus will be the development of English skills: listening comprehension, reading, writing, and speaking. A key objective will be to increase and broaden knowledge of technical and scientific vocabulary. The final result; students should be able to produce English presentations, especially science-related presentations. The emphasis of the course content is intended to strengthen and expand both comprehension and expressive abilities. It is anticipated that this will include improving the basic aspects of clear writing: basic grammar, vocabulary, and recurring patterns of typical sentence structures.

【キーワード】 Technical English，テクノロジー， English presentations 【先行科目】『初級技術英語』 (1.0) 【関連科目】『上級技術英語』 (1.0) 【履修要件】 You should have received credit for ”Basic Technical English.” 【履修上の注意】 This subject is a compulsory for Nichia STC students. If you are not

a Nichia STC student and want to attend this class, please contact a coordinator in your department for getting permission for attendance.

【到達目標】

1. To develop vocabulary, sentence formations, and presentation principles of communication in order to establish a firm foundation for public speaking and technical presentations, in English.

2. To develop English skills: listening comprehension, speaking, reading and

writing. 【授業計画】

1. Conventions of English 1

2. Conventions of English 2

3. Vocabulary: correct spelling

4. Vocabulary: how to communicate to an English-speaking person the spelling of a word or phrase, homonyms, synonyms, acronyms, and commonly misspelled or misunderstood English words.

5. Noun and verb agreement 1

6. Noun and verb agreement 2 7. Reading skills 1

8. Reading skills 2

9. Various types of Writing 1

10. Various types of writing 2

11. Important points in public speaking 1

12. Important points in public speaking 2

13. How to describe something or someone: commonly used adjectives, phrases and words that describe location. phrases and words that describe the frequency of occurrence or the unique qualities of a thing, a process, or research

14. How to compose communication: letters of acceptance, letters of inquiry, for example.

15. Electronic forms of communication (E-mail), telephone conversations, how to cite internet sources of information, plagiarism.

【成績評価基準】 Grades will be based on class participation, a mid-term, presentations

and homework assignments. 【学習教育目標との関連】 [JABEE 関連] 【教科書】”Presenting Science, ” (Second Edition), 2008, Macmillan Language-

House. 【参考書】 [参考資料] 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216144

【対象学生】 Nichia STC students and, possibly, a limited number of other students

are permitted to take this course. 【連絡先】

⇒ カーペンター (共通講義棟 1F 国際連携教育開発センター, 088-656-7643, walter@cicee.tokushima-u.ac.jp) ����

Technical English 1 unit (selection)

Walter Carpenter · Designated Associate Professor / Production Systems Engineering, Department of Mechanical Engineering

Target⟩ The aim of the course is to enhance reading, writing, speaking and listening skills. Another goal is to establish a firm foundation for students to present their knowledge to others, in English.

Outline⟩ The primary focus will be the development of English skills: listening comprehension, reading, writing, and speaking. A key objective will be to increase and broaden knowledge of technical and scientific vocabulary. The final result; students should be able to produce English presentations, especially science-related presentations. The emphasis of the course content is intended to strengthen and expand both comprehension and expressive abilities. It is anticipated that this will include improving the basic aspects of clear writing: basic grammar, vocabulary, and recurring patterns of typical sentence structures.

�������⟩ Technical English, technology, English presentations Fundamental Lecture⟩ “Basic Technical English”(1.0) Relational Lecture⟩ “Advanced Technical English”(1.0) Requirement⟩ You should have received credit for ”Basic Technical English.” Notice⟩ This subject is a compulsory for Nichia STC students. If you are not a

Nichia STC student and want to attend this class, please contact a coordinator in your department for getting permission for attendance.

Goal⟩

1. To develop vocabulary, sentence formations, and presentation principles of communication in order to establish a firm foundation for public speaking and technical presentations, in English.

2. To develop English skills: listening comprehension, speaking, reading and

writing.

Schedule⟩

1. Conventions of English 1

2. Conventions of English 2

3. Vocabulary: correct spelling

4. Vocabulary: how to communicate to an English-speaking person the spelling of a word or phrase, homonyms, synonyms, acronyms, and commonly misspelled or misunderstood English words.

5. Noun and verb agreement 1

6. Noun and verb agreement 2 7. Reading skills 1

8. Reading skills 2

9. Various types of Writing 1

10. Various types of writing 2

11. Important points in public speaking 1

12. Important points in public speaking 2

13. How to describe something or someone: commonly used adjectives, phrases and words that describe location. phrases and words that describe the frequency of occurrence or the unique qualities of a thing, a process, or research

14. How to compose communication: letters of acceptance, letters of inquiry, for example.

15. Electronic forms of communication (E-mail), telephone conversations, how to cite internet sources of information, plagiarism.

Evaluation Criteria⟩ Grades will be based on class participation, a mid-term,

presentations and homework assignments. Relation to Goal⟩ [JABEE 関連] Textbook⟩ ”Presenting Science, ” (Second Edition), 2008, Macmillan Language-

House. Reference⟩ [参考資料] Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216144

Student⟩ Nichia STC students and, possibly, a limited number of other students

are permitted to take this course.

Contact⟩

⇒ Carpenter (共通講義棟 1F 国際連携教育開発センター, +81-88-656-7643, walter@cicee.tokushima-u.ac.jp) ����